A learner mass(75 kg) is standing next to the road and sees a stationary car (mass 25kgl in front of her.she decides to run towards the car and jump on it and roll down the street.The learner accelerates at a constant acceleration of 2m/s calculate the car'a velocity just after the learner jumped on it

1
Expert's answer
2022-03-20T18:49:54-0400

"m_1v=(m_1+m_2)u,"

"u=v\frac{m_1}{m_1+m_2}=1.5~\frac ms."